OpenStudy (anonymous):

\[ \frac{d}{d(x^2)}x \]The derivative of \(x\) with respect to \(x^2\). I'm getting \(1/(2x)\) but wolfram is getting just \(1/x\). I wonder how they are getting that.

OpenStudy (nincompoop):

how are you getting that?

OpenStudy (anonymous):

In the general case: \[ \frac{d}{d(x^2)}x^n \]I would say this is \[ \frac{n}{2}x^{n-2} \]But wolfram does not get that \(n/2\) coefficient.

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

@nincompoop I used substitution.

OpenStudy (nincompoop):

I know, link the wolfram

OpenStudy (nincompoop):

Simplify the following: (d x)/(d x^2) Cancel common terms in the numerator and denominator of (d x)/(d x^2). (d x)/(d x^2) = d/d×x/x^2 = x/x^2: x/x^2 For all exponents, a^n/a^m = a^(n-m). Apply this to x/x^2. Combine powers. x/x^2 = x^(1-2): x^1-2 Evaluate 1-2. 1-2 = -1: Answer: x^-1

OpenStudy (nincompoop):

that d is not coming out as derivative operation

OpenStudy (anonymous):

Here is one way I considered doing it: \[ \frac{d}{d(x^2)}x = \frac{d}{2xdx}x = \frac{1}{2x}\frac{d}{dx}x = \frac{1}{2x} \]

OpenStudy (anonymous):

Another way: \(u = x^2\) and \(x = u^{1/2}\): \[ \frac{d}{d(x^2)}x = \frac{d}{du}x = \frac{d}{du}u^{1/2} = \frac{1}{2u^{1/2}} = \frac{1}{2x} \]

OpenStudy (nincompoop):

http://www.wolframalpha.com/input/?i=differentiate+x+respect+to+x%5E2 no evaluation

OpenStudy (anonymous):

Suppose \[ \frac{d}{d(x^2)}x = y \]Then \[ dx = yd(x^2)=y(2x)dx \implies 1 = y(2x)\implies y = \frac{1}{2x} \]

OpenStudy (anonymous):

Is wolfram pelletting the bed?

OpenStudy (nincompoop):

laughing out loud I am trying to get the right command http://www.wolframalpha.com/input/?i=dy%2Fdx%5E2+x+

OpenStudy (anonymous):

What does wolfram think it is doing when it gets \(1/x\)...

OpenStudy (anonymous):

or \(x^{n-2}\) for that matter.

OpenStudy (agent0smith):

Hmm... http://www.wolframalpha.com/input/?i=d%2F%5Bd%28x%5E2%29%5D+x%5E4and x^3 does the same, there's no constant out the front.

ganeshie8 (ganeshie8):